Operations Report – March 30th

Crew 208 Operations Report 30-MAR-2019

SOL: 2

Name of person filing report: Amanda Manget

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on the generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): 15

Solar – SOC 99% @ 1900 (before David turned on the generator for the night)

Diesel Reading – 48%

Propane Reading – 56%

Ethanol Free Gasoline – 2 gallons

Water (auxiliary tank) – 0 gallons (Note: Not in Use)

Water (static tank) – 228 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: 0

Water in GreenHab – 114 gallons

Static to Loft Pump used – Yes

Water Meter: 142228.4 gallons @ 2000

Toilet tank emptied: No

Deimos rover used: No, not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: No

Hours: 85.1

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Opportunity rover used: No

Hours: 54.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: No

Hours: 84.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Not Used

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) N/A

Reason for use: Not Used

Oil Added? No

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: None

HabCar used and why, where?

Hab car was driven to Hanksville to pick up water for Crew 208

Har car was driven to Grand Junction, CO to pick up 2 members of Crew 208

CrewCar used and why, where? Not Used

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: EV suits normal, all batteries read >12.5 volts

Summary of Hab operations: Checked SOC of the electrical system as follows:

0930 SOC 100%, generator turned off by David
1500 SOC 100%, generator off
1900 SOC 99%, generator turned on by David

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: None (Note: observatory is closed)

Summary of health and safety issues: None

Greenhab Report – March 30th

Crew 208 GreenHab Report – 30-March-2019

GreenHab Officer: Erika Rydberg

Environmental control:

Ambient with the door open

Shade cloth (80%) on

Average temperatures: 22 degrees Celsius

Low: 12
High: 29

Hours of supplemental light: 0

Daily water usage for crops: 6 gallons

Water in Blue Tank — 114 gallons

Time(s) of watering for crops:10:45 and 16:55

Changes to crops: None

Narrative: First day for me in the Greenhab, I garden at home, so it was nice to see some good growth coming from vegetables in this environment. Herbs are abundant.

Harvest: 2 grams basil

Support/supplies needed: None

Sol Summary – March 30th

Crew 208 Sol Summary Report 30-MAR-2019

Sol: 2

Summary Title: Welcome to MDRS, Crew 208!

Author’s name: Julielynn Wong

Mission Status: Nominal

Sol Activity Summary:

GHO arrived yesterday PM.

COM trained GHO on filling out 3D print logs.

Crew 208 completed 3D printing in the Hab today for the thoracentesis training study.

XO continued to process MDRS maps from drone images. GHO prepared 3D print files of our drone maps and started printing our 1st 3D map of MDRS. XO + GHO conducted several drone flights today.

XO + David traveled to Hanksville to refill our water resources at MDRS.

COM trained incoming GHO on GreenHab duties and reports. XO trained incoming ENG on reporting duties.

David drove CrewCar to GJT to meet Crew 208 ENG + HSO + transport to MDRS.

Crew 208 will drive into Hanksville tonight for dinner after reports have been submitted to CAPCOM.

Look Ahead Plan: Continue 3D printing maps of MDRS in Hab. MDRS orientation and training to commence on Sol 2 and finish on Sol 3. COM to give a tour to PS students from 1400-1600 on Sol 3. Begin sim in PM of Sol 3.

Anomalies in work: None

Weather: Sunny

Crew Physical Status: Nominal

EVA: n/a

Reports to be filed: Operations, Greenhab

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site